OpenStudy (anonymous):

Lance has a moment of inertia of 420 kg m2 while riding the Rotor. The floor of the Rotor drops out from under its riders once it reaches the angular velocity (ω omega) of 1.6 radians per second. Calculate his angular momentum. Remember to identify all data (givens and unknowns), list equations used, show all your work, include units, and the proper number of significant digits to receive full credit. @Abhisar @ganeshie8 @Nnesha @zepdrix @TheSmartOne @pooja195

OpenStudy (anonymous):

PLEAAAAAAAAAASEEEEE HEEEELPP!!

pooja195 (pooja195):

1.6*pi radians / second is the angular speed of the rotor.

pooja195 (pooja195):

Do you know that much?

OpenStudy (anonymous):

no i really need help with it

pooja195 (pooja195):

Given the moment of inertia or 420 kg m^2, multiply it by the angular speed to get Lance's angular momentum.

pooja195 (pooja195):

Try it out :)

OpenStudy (anonymous):

is that all i need to do?

pooja195 (pooja195):

I think so i dont see anything else that you would need to do unless im doing this wrong =/

pooja195 (pooja195):

@shamim ?

OpenStudy (anonymous):

wat do u think @shamim ?

pooja195 (pooja195):

I would think thats all you need to do

OpenStudy (anonymous):

ok

OpenStudy (shamim):

Ok. Do u know p=mv

OpenStudy (anonymous):

no

OpenStudy (shamim):

p=linesr momentum

OpenStudy (shamim):

m=mass of that body

OpenStudy (anonymous):

ikd any of this.. but anything that'll get this problem slved will help

OpenStudy (shamim):

v=velocity of that object

OpenStudy (anonymous):

so wat do i need to do i dont have all the time.

pooja195 (pooja195):

is this a test?

OpenStudy (anonymous):

no! why do u always ask me that question! its my physics hw if it was a test I would've never posted it on here.

pooja195 (pooja195):

Was just curious anyways continue on you said im running out of time

OpenStudy (anonymous):

@TheSmartOne HELP MEEEE ANYONE PLZZZ!!!!

OpenStudy (shamim):

P=mv..... for linear motion L=I*omega.....for angular motion

OpenStudy (anonymous):

ok?

OpenStudy (shamim):

Now for ur math problem u need to use the formuls L=I*omega

OpenStudy (shamim):

Here L=angular momentum

OpenStudy (shamim):

I=moment of inertia

OpenStudy (anonymous):

so basically wat pooja suggested?..

OpenStudy (shamim):

Omega=angular velocity

OpenStudy (anonymous):

so basically wat pooja suggested?..

OpenStudy (shamim):

Given I=420kg.m^2 Omega=1.6rad/sec

OpenStudy (shamim):

L=?

TheSmartOne (thesmartone):

\[\sf Angular~Momentum = Moment ~ of ~ inertia \times Angular~Velocity\] And in symbols that would be \[\sf L = I \times \omega \]

TheSmartOne (thesmartone):

They have given us that \(\sf Moment~ of~inertia= 420~kg m^2\)

OpenStudy (anonymous):

right

TheSmartOne (thesmartone):

And that \(\sf Angular~velocity = 1.6~radians~per~ second\)

OpenStudy (anonymous):

correct :)

TheSmartOne (thesmartone):

And since it is in radians we don't need to change it to degrees because it has to be in radians :)

OpenStudy (anonymous):

okay!

TheSmartOne (thesmartone):

Now we just plug it in :)

TheSmartOne (thesmartone):

\(\Large\sf Angular~momentum= 420 ~kg~m^2\times 1.6 \frac{rad}{sec}\)

OpenStudy (anonymous):

Oh wooow! okaayyy!

OpenStudy (anonymous):

is it okay if you can help me answer this or explain this in complete sentences plz?

OpenStudy (anonymous):

:)

TheSmartOne (thesmartone):

So basically, you put the formula we used at the top. And then write what we were given. And then show how we plugged it in. And how we solved for it and got the correct units. And then round to the correct number of significant figures.

TheSmartOne (thesmartone):

And ta-da you receive full credits :)

OpenStudy (anonymous):

Okay I will work on that and have you check if it looks right for me plz :)

TheSmartOne (thesmartone):

Sure :)

TheSmartOne (thesmartone):

I honestly thought this problem was hard, until I remembered the formula xD All those sentences confused me. I had to remove all that useless info and just take the numbers they gave me. Haha

OpenStudy (anonymous):

hahaha that is true the sentences really got me at first, but then with the formula and when you showed me how to do it, it was a piece of cake!

OpenStudy (anonymous):

i have a question @pooja195

OpenStudy (anonymous):

I can clearly see that you are a barbie girl. Oh baby baby..

OpenStudy (anonymous):

Awk moment when this is from the physics test

TheSmartOne (thesmartone):

o.o